Decide whether each of these integers is congruent to 3 modulo 7.
(a) 37
(b) 66
(c) -17
(d) -67
1
Expert's answer
2020-10-19T17:51:06-0400

Solution. If a and b any integer and m is a positive integer then a congruent to b modulo m when m is divides the difference a-b. Therefore

(a)


"\frac{37-3}{7}=\frac{34}{7}"

37 is not congruent to 3 modulo 7.

(b)


"\frac{66-3}{7}=\frac{63}{7}=9"

66 is congruent to 3 modulo 7.

(c)


"\frac{-17-3}{7}=\frac{-20}{7}"

-17 is not congruent to 3 modulo 7.

(d)


"\frac{-67-3}{7}=\frac{-70}{7}=-10"

-67 is congruent to 3 modulo 7.

Answer. (a) 37 is not congruent to 3 modulo 7; (b) 66 is congruent to 3 modulo 7; (c) -17 is not congruent to 3 modulo 7; (d) -67 is congruent to 3 modulo 7.